OpenStudy (anonymous):

solve by u-substitution x^6-9x^3=-8

OpenStudy (anonymous):

let u = x^3 therefor the equation turns to u^2-9u=-8 u^2-9u+8=0 (u-8)(u-1)=0 now we have u=8 or u=1 now you can put x^3 back into the u position x^3=8 or x^3=1 then you can solve the problem from there :) let me know if that made sense or not

OpenStudy (jdoe0001):

\(\bf {\color{blue}{ x^6\implies (x^3)^2\qquad \qquad u=x^3}} \\ \quad \\ \quad \\ x^6-9x^3=-8\implies (x^3)^2-9x^3+8=0\implies {\color{blue}{ u}}^2-9{\color{blue}{ u}}+8=0\)

OpenStudy (anonymous):

Thank you so much, i was stuck at how to make it equal to zero, but now i can solve from here. Thanks!

OpenStudy (jdoe0001):

yw

OpenStudy (tkhunny):

And then, of course, keep factoring. You're nowhere near done!

OpenStudy (anonymous):

Glad to help!!!